Given: TTGGGTTAGGGTTAGGGTTAGGA

What is the probability that the specific sequence required to form a protein structure occurs randomly in any given DNA? Assume that all four DNA nucleotide bases (A, T, C and G) are equally probable and nucleotides are present in excess.
Answer #1

Probability of any DNA base = 1/4 = 0.25

Now the sequence is = TTGGGTTAGGGTTAGGGTTAGGA

Total DNA base = 23

Probability of any specific 23 base DNA will be = (1/4)23    = (0.25)23 =  1.42109E-14